### v1

# Rethink the Evaluation Protocol of Model Merging on Classification Task

Abstract
Model merging combines multiple fine-tuned models into a single one via parameter fusion, achieving improvements across many tasks. However, in the classification task, we find a misalignment issue between merging outputs and the fine-tuned classifier, which limits its effectiveness. In this paper, we demonstrate the following observations: (1) The embedding quality of the merging outputs is already very high, and the primary reason for the differences in classification performance lies in the misalignment issue. (2) We propose FT-Classifier, a new protocol that fine-tunes an aligned classifier with few-shot samples to alleviate misalignment, enabling better evaluation of merging outputs and improved classification performance. (3) The misalignment is relatively straightforward and can be formulated as an orthogonal transformation. Experiments demonstrate the existence of misalignment and the effectiveness of our FT-Classifier evaluation protocol.
Rethink the Evaluation Protocol of Model Merging on Classification Task
[…]
1 Introduction
Fine-tuning a pre-trained model (PLM) on downstream tasks to obtain a task-specific fine-tuned model is a common practice nowadays Devlin et al. (2018). Building on this training paradigm, real-world applications increasingly aim to leverage a single model that can be fine-tuned across multiple tasks, thus enabling it to address all tasks simultaneously Zhang and Yang (2021). This is known as multi-task learning (MTL), which generally aggregates all task data for joint training. However, MTL not only requires the full amount of task data, which raises data privacy concerns, but also involves a fully fine-tuning process, leading to high training costs and poor scalability Yang et al. (2024b). Therefore, researchers are actively exploring model merging as an alternative Izmailov et al. (2018); Jin et al. (2022); Ilharco et al. (2022); Yang et al. (2024c, a).
Model merging aims to directly combine multiple homologous fine-tuned task-specific models in the parameter space, without requiring the full dataset and the fine-tuning process, to obtain a merged model capable of solving multiple tasks simultaneously Yadav et al. (2024). Several studies have investigated model merging in areas such as continual learning Marczak et al. (2024), information retrieval Li et al. (2024), and classification Yu et al. (2024). Interestingly, we find that compared to other tasks, model merging brings fewer benefits in classification tasks.
We argue that this is because classification tasks typically require a classifier (e.g., the final layer of a neural network) for evaluation. However, since the classifier’s weight dimensions vary across tasks (due to different numbers of classes), classifiers cannot be merged directly. As a result, evaluation of the merged model depends on the fine-tuned classifier, causing a misalignment between the merging output and the classifier, which ultimately leads to suboptimal performance in classification tasks.
To support our argument, we visualize the merging outputs of three representative methods on text classification datasets using T-SNE Van der Maaten and Hinton (2008), as shown in Figure 1. From the figure, we observe two key phenomena:
Refer to caption Refer to caption
(a) Weight Averaging (79.2%) (b) Weight Averaging (55.1%)
Refer to caption Refer to caption
(c) Task Arithmetic (82.9%) (d) Task Arithmetic (81.5%)
Refer to caption Refer to caption
(e) Ties (83.3%) (f) Ties (83.4%)
Figure 1: T-SNE visualization of three representative merging methods on AG News and DBpedia. The blue points represent the sentence embeddings of the fine-tuned model, while the red points represent the results of the merging model. The color shades indicate different labels. The value indicates the classification accuracy of merging methods. For results on other datasets, please refer to Appendix A.1.
(1) Representation bias between merging and fine-tuned models. The overall red and blue clusters are non-overlapping, indicating that merging outputs and fine-tuned model outputs exhibit a bias in the latent space. This phenomenon is consistent with the observations in Yang et al. (2024b).
(2) Good clustering effects in merging outputs. Despite notable gaps in classification accuracy between different merging methods (e.g., nearly a $30\%$ accuracy difference between Weight Averaging and Ties on DBpedia), their merging outputs exhibit similar clustering performance. This suggests that merging outputs retain high-quality embeddings, comparable to fine-tuned models.
These observations lead us to suspect reasonably: There is a representation misalignment issue between merging outputs and the fine-tuned classifier, and the current evaluation protocol appears incapable of accurately assessing the true quality of merging methods. Therefore, in this work, we conduct the following analysis to validate our suspicion.
Assessing the quality of merging outputs
We bypass the classifier and instead use a few-shot set of labeled samples as class anchors. Employing a KNN-based method—commonly applied in sentence representation studies—we directly measure classification precision. The results are reported in Figure 2. Surprisingly, the KNN-based evaluation outperforms the current evaluation protocol on three representative methods, even with limited labeled data, demonstrating the high quality of the merging outputs. Moreover, this also indicates that the performance drop of the current evaluation protocol mainly contributes to the misalignment issue.
Alignment formulation
We designed two fine-tuning strategies: (1) adding a mapping matrix between the merging output and the fine-tuned classifier, and (2) directly fine-tuning the classifier to adapt to the merging output. The results are shown in Figure 4. The findings confirm that aligning either the merging output or the classifier can significantly improve classification performance, with results even approaching the upper limit (the fine-tuned model’s performance).
New evaluation protocol for misalignment alleviation
Based on the above observations, we believe that the current evaluation protocol mainly assesses the compatibility between the merging output and the fine-tuned classifier rather than effectively evaluating the quality of the merging outputs themselves, which in turn limits the final classification performance. Therefore, we propose the FT-Classifier evaluation protocol, which leverages few-shot unlabeled data and employs minimal training steps without introducing additional parameters, aiming to train an aligned classifier and then make the evaluation more convincing. The results of the FT-Classifier evaluation protocol, presented in Figure 5, demonstrate that only a minimal amount of data is required for the alignment, resulting in a significant improvement in classification performance.
The misalignment is an orthogonal transformation
The fact that the misalignment issue can be mitigated with only a small amount of data demonstrates that the transformation of misalignment is exceedingly straightforward. Intuitively, we hypothesize that this transformation can be approximated as a sufficiently simple orthogonal transformation. To support this, we add an orthogonal transformation regularization to the mapping matrix. The results shown in Table 2 represent that it is possible to train an orthogonal transformation with a slight drop in classification performance. This means that, through a rigid transformation such as rotation or reflection, the merging outputs can be aligned with the fine-tuned classifier. It indicates that the merging outputs, compared to the fine-tuned model’s outputs, preserve the intrinsic properties of the manifold, which are essential for classification.
In summary, we make the following contributions:
  • Through empirical analysis, we reveal that the misalignment issue between merging outputs and the fine-tuned classifier renders the current evaluation protocol ineffective for accurately assessing merging methods.
  • We introduce a new FT-Classifier evaluation protocol that alleviates the misalignment issue by limited data, enabling a more accurate assessment of merging methods while significantly enhancing their performance.
  • We prove that this misalignment can converge to an orthogonal transformation, indicating that the intrinsic properties of the manifolds in the merging outputs remain consistent with those of the fine-tuned model.
2 Background
2.1 Definition of Model Merging
Let $f_{b}$ denote the pre-trained base model with parameter $\mathbf{\Theta}_{b}$. Given $T$ specific tasks with training data $X_{t}$ and corresponding label $Y_{t}$ for each, one can obtain $T$ fine-tuned models. And the fine-tuned model for task $t$ can be denoted as $f_{t}$ with parameter $\mathbf{\Theta}_{t}$. Without fully fine-tuning, model merging aims to merge these homologous models to generate a single $f_{m}$ with parameter $\mathbf{\Theta}_{m}$, which can simultaneously perform well across all $T$ tasks.
2.2 Representative Methods
Up to date, numerous model merging approaches have been proposed, which we will discuss in detail in Section 7. In this section, we primarily focus on introducing the three most representative methods. Weight averaging Wortsman et al. (2022) simply averages $T$ fine-tuned models for merging, which is formulated as:
$\mathbf{\Theta}_{m}=\frac{1}{T}\sum_{t=1}^{T}\mathbf{\Theta}_{t}.$ (1)
Task Arithmetic Ilharco et al. (2022) defines the difference between the fine-tuned model and the pre-trained model in parameter space as a task vector, denoted by $\mathbf{\Delta}_{t}$. Experiments demonstrate that performing arithmetic operations on task vectors can effectively guide the merged model. The process is defined as:
$\mathbf{\Delta}_{t}$ $=\mathbf{\Theta}_{t}-\mathbf{\Theta}_{b},$ (2)
$\mathbf{\Theta}_{m}$ $=\mathbf{\Theta}_{b}+\lambda\sum_{t=1}^{T}\mathbf{\Delta}_{t},$
where $\lambda\in[0,1]$ is a scaling term determined by held-out validation sets.
Ties Yadav et al. (2024) posits that parameter conflicts are widespread across multiple task vectors. To address this, it preserves task parameters with large magnitudes and resolves sign conflicts using an election mechanism. We denote the task vector processed by Ties as $\mathbf{\Delta}_{t}^{\rm Ties}$, then the parameters merged by Ties are calculated as follows:
$\mathbf{\Theta}_{m}=\mathbf{\Theta}_{b}+\lambda\sum_{t=1}^{T}\mathbf{\Delta}_{t}^{\rm Ties}.$ (3)
2.3 Current Evaluation Protocol
Existing merging methods primarily focus on generating a comprehensive encoder that absorbs the ability from each fine-tuned model. While the merging encoders $f_{m}$ are evaluated on the fine-tuned classier, which can be formulated as:
$\hat{Y}_{t}=f_{m}(X_{t};\mathbf{\Theta}_{m})\cdot\mathbf{W}_{t},$ (4)
where $\mathbf{W}_{t}$ represents the weight of the classifier of task $t$. It maps the merging outputs to the predicted labels $\hat{Y}_{t}$.
3 Existence of the Misalignment Issue
From Eq. 4, we can easily find the fact that due to the change in parameters, the merging outputs of $X_{t}$ generated by $f_{m}$ are inevitably shifted in the latent space compared to those generated by $f_{t}$. However, the evaluation of $f_{m}$ still relies on $\mathbf{W}_{t}$, which is tailored to $f_{t}$. Naturally, this raises a question: Does the outputs of $f_{m}$ align well with the fine-tuned classifier $\mathbf{W}_{t}$?
3.1 Merging outputs Quality Evaluation
As shown and analyzed in Figure 1, we speculate that the merging outputs produced by different $f_{m}$ consistently exhibit high quality. To validate our hypothesis, we use a KNN-based method to bypass the classifier and directly evaluate the classification performance of different merging outputs. Illustrations of the current evaluation and this KNN-based one are presented in Figure 3 (a) and (b).
Specially, for the KNN evaluation, we first randomly select few-shot labeled data as anchors for each class, then calculate the average Euclidean Distance between the test sample and these anchors for each class, and finally assign the class with the minimum distance as the prediction. The overall process is formulated as:
$\hat{y}_{t}=\arg\min_{y_{t}\in Y_{t}}\frac{1}{k}\sum_{i=1}^{k}d\left(X_{t},[A_{t}^{y_{t}}]^{i}\right).$ (5)
where $y_{t}$ is an element of $Y_{t}$, $k$ refers to the number of few-shot samples, $d$ represents the Euclidean Distance metric, $A_{t}^{y_{t}}$ represents the randomly selected anchors for label $y_{t}$, and []delimited-[][$\cdot$][ ⋅ ] denotes the indexing function. From Eq. 5, we observe that the KNN evaluation directly utilizes merging outputs for classification with only few-shot labeled data, and we believe this approach provides an effective assessment of merging outputs quality.
Refer to caption Refer to caption Refer to caption
Figure 2: The performance of KNN evaluation compared with current evaluation protocol under varying numbers of few-shot examples $k$. The $x$-axis represents the $k$, and the $y$-axis represents the average of Accuracy for all merging tasks. Experimental details and explanations are provided in Section 6.
Results of KNN evaluation compared with the current evaluation protocol and upper limits fine-tuned models are shown in Figure 2. From the figure, we observe that directly using the KNN evaluation yields excellent classification results. Even with only 5 few-shot examples, the KNN evaluation outperforms the current protocol, which relies on a fine-tuned classifier. This suggests that the outputs of different merging methods are of high quality, and the poor performance in the current evaluation protocol is more likely caused by the misalignment between the merging outputs and the fine-tuned classifier.
3.2 Revisiting the Objective of Model Merging
Revisiting the target of model merging, which aims to generate a new model $f_{m}$ using $f_{1},...,f_{T}$ that performs well across all $T$ tasks, we note that the upper bound of $f_{m}$’s performance on a specific task $t$ corresponds to the result from fine-tuned model $f_{t}$. Naturally, the objective of model merging is to minimize this performance discrepancy of $f_{m}$ and $f_{t}$ Ortiz-Jimenez et al. (2024), which is defined as:
$\arg\min_{\mathbf{\Theta}_{m}}\sum_{t=1}^{T}g(f_{m}(X_{t};\mathbf{\Theta}_{m})\cdot\mathbf{W}_{t},f_{t}(X_{t};\mathbf{\Theta}_{t})\cdot\mathbf{W}_{t})$ (6)
where $\cdot$ represents matrix multiplication, and $g$ is a measurement function.
From the objective, we observe that the optimal $\mathbf{\Theta}_{m}^{*}$ should not only be related to the fine-tuned model $\mathbf{\Theta}_{t}$ but also to $\mathbf{W}_{t}$. Since we have also observed that the merging output and classifier exhibit a certain misalignment, we hypothesize that alleviating the misalignment in latent space could lead to a better evaluation of $f_{m}$.
Figure 3: Illustrations of some variants referred to in this paper.
3.3 Alignment Formulation
Consequently, we conduct two simple strategies to align merging outputs and the fine-tuned classifier: (1) introducing a mapping function between the outputs and the classifier to learn the alignment, and (2) directly fine-tuning the classifier to better align with the outputs. Both training approaches used Eq. 6 as the objective, with KL divergence Kullback and Leibler (1951) as the metric function. Following Yang et al. (2024b), we employ unlabeled test data to train the related parameters.
Specifically, for the first approach, we simply define the mapping function as a square matrix $\mathbf{M}_{t}$ with the same dimensionality as the hidden layer, and the optimization objective, named Aligned $\mathbf{M}_{t}$ is defined as follows:
[…]
Note that during the training process, only $\mathbf{M}_{t}$ is updated, while all other parameters remain frozen.
For the second approach, we use $\mathbf{W}_{t}$ to initialize a new classifier $\mathbf{W}_{t}^{\prime}$, and define the optimization objective Aligned $\mathbf{W}_{t}^{\prime}$ as follows:
$\arg\min_{\mathbf{W}_{t}^{\prime}}\sum_{t=1}^{T}{\rm KL}(f_{m}(X_{t};\mathbf{\Theta}_{m})\!\cdot\!\mathbf{W}_{t}^{\prime},f_{t}(X_{t};\mathbf{\Theta}_{t})\!\cdot\!\mathbf{W}_{t})$ (8)
In this objective, only the new classifier $\mathbf{W}_{t}^{\prime}$ requires updating.
Figure 4: Comparison of the Current Eval and two Aligned variants.
Illustrations of Aligned $\mathbf{M}_{t}$ and Aligned $\mathbf{W}_{t}^{\prime}$ are shown in Figure 3 (c), (d), respectively. The relevant comparable results are shown in Figure 4. From the Figure, we observe that:
(1) With only a limited number of parameters, the misalignment can be captured, achieving surprising classification performance approaching the upper bound of the fine-tuned model. Both aligned variants require very few parameters, with the Aligned $\mathbf{W}_{t}^{\prime}$ variant even using the exact same number of parameters as the fine-tuned classifier $\mathbf{W}_{t}$. This also indicates that the quality of outputs via different merging methods is already high.
(2) The quality of the outputs produced by different $f_{m}$ does not differ significantly, and the current evaluation protocol does not assess $f_{m}$ well. Since the performance gap of different aligned variants is much smaller than the current evaluation of $f_{m}$. For example, the performance of Weight Averaging was initially around 10% lower than that of Task Arithmetic, but after alignment, their performances became nearly identical.
These findings demonstrate that the misalignment issue hinders the current evaluation protocol from effectively assessing the capacity of $f_{m}$. And alleviating the misalignment problem could potentially improve classification performance.
Method AG News Yelp Amazon Yahoo DBpedia Avg.
FT 91.6 60.4 57.9 72.7 98.8 76.3
MTL 90.6 59.1 55.6 71.3 98.5 75.0
WA 79.2 49.8 45.0 50.3 55.1 55.8
+FT-C 85.9 49.8 46.8 66.1 96.5 69.0
TA 82.9 55.8 48.4 53.1 81.5 64.3
+FT-C 87.5 57.0 52.5 67.9 96.6 72.3
Ties 83.3 55.7 47.0 55.9 83.4 65.1
+FT-C 87.5 58.2 52.2 69.1 97.7 73.0
Table 1: Comparison of FT-Classifier evaluation with the corresponding current one. The results in this Table are based on a few-shot setting of $k=5$. FT represents the results of the Fine-tuned model, MTL represents the results of traditional Multi-task Learning, WA represents Weight Averaging, TA represents Task Arithmetic, and +FT-C represents +FT-Classier evaluation protocol.
3.4 Further Discussion
Model merging involves combining the parameters of multiple $f_{t}$ models to generate $f_{m}$. Consequently, the output of $f_{m}$ is inevitably shifted in the latent space compared to the original $f_{t}$. This shift is unavoidable. In this case, using $\mathbf{W}_{t}$, which is aligned well with $f_{t}$, to evaluate $f_{m}$ is inappropriate. The evaluation result reflects not the capability of $f_{m}$, but rather its compatibility with the fine-tuned classifier $\mathbf{W}_{t}$. Moreover, through experiments, we find that simply aligning the merging outputs and the classifier in the latent space can significantly improve the classification performance of $f_{m}$. Naturally, this motivates us to propose a new and reasonable evaluation protocol for model merging. This protocol would not only enable a more precise assessment of $f_{m}$, but also enhance classification performance through mitigating the misalignment issue.
4 FT-Classifier Evaluation Protocol for Alleviating the Misalignment Issue
In model merging, obtaining few-shot data for each task is a relatively flexible requirement. Many studies use few-shot data to guide the merging methods Xiao et al. (2023), and some even rely on large-scale labeled or unlabeled data to determine merging weights Yang et al. (2024c); Jin et al. (2022) or the parameters to be pruned Ortiz-Jimenez et al. (2024); Zhu et al. (2024). In our proposed evaluation protocols, we also incorporate few-shot samples for misalignment alleviation. Meanwhile, as the goal of model merging is to produce a single model capable of flexibly addressing various tasks, we argue that the evaluation protocol should avoid introducing additional parameters. Hence, we select the Aligned $\mathbf{W_{t}^{\prime}}$ variants to mitigate the misalignment.
In particular, we adopt few-shot samples to fine-tune an aligned classifier for merging model evaluation. Compared to the current evaluation protocol, this approach introduces an additional post-processing step, while keeping the parameter scale and network structure unchanged. Since only a few-shot samples are used for training, the required time remains entirely acceptable. The training objective of this approach is consistent with Eq. 8, also called FT-classifier Evaluation, with the only difference being the use of few-shot data.
Comparable results of the current evaluation and FT-Classifier are presented in Figure 5 and Table 1. From the Figure, we observe that after simple fine-tuning even on 1-shot samples, FT-Classifier Eval performs significantly better than the Current Eval. This result leads us to wonder whether the alignment is truly straightforward; otherwise, it should not have been captured so easily.
Refer to caption Refer to caption Refer to caption
Figure 5: The performance of different evaluation protocols under varying numbers of few-shot examples $k$. Corresponding detailed results while $k=5$ are shown in Table 1.
5 The Misalignment is an Orthogonal Transformation
Intuitively, we hypothesize that this transformation is the simplest form of an orthogonal matrix, meaning that the merging outputs can be aligned in the latent space through simple rotations or reflections. To accomplish this, we enforce $\mathbf{M}_{t}$ to be an orthogonal matrix by adding a regularization constraint while training. Specifically, we introduce a regularization loss on Aligned $\mathbf{M}_{t}$ to ensure this property. The objective function is defined as follows:
[…]
where $\alpha$ is the weight hyperparameter, and $\mathbf{I}$ is an identity matrix.
Method Unaligned + Orth $\mathbf{M}_{t}$ + $\mathbf{M}_{t}$
WA 55.8 72.8 74.9
TA 64.3 72.7 74.5
Ties 65.1 73.8 74.7
Table 2: Comparison of realigned current eval, orthogonal transformation alignment (+Orth $\mathbf{M}_{t}$), and normal alignment (+ $\mathbf{M}_{t}$).
Detailed results of the orthogonal regularization are shown in Table 2. From the Table, we observe that when $\mathbf{M}_{t}$ converges to an orthogonal transformation, the final classification performance is not significantly affected (only nearly 2% dropped). This demonstrates that through an orthogonal transformation, the merging outputs can be aligned with the fine-tuned classifier. Since orthogonal transformations in space are rigid transformations, such as rotations or reflections, this suggests that these merging methods preserve the essential capabilities of the fine-tuned models, while the intrinsic properties of the manifold in the latent space remain unchanged.
6 Experiments
The above observations demonstrate that alleviating the misalignment and modifying the evaluation protocol is necessary. And our new evaluation protocols more accurately assess the performance of merging methods and are able to improve the overall classification performance. In this Section, we will provide more detailed analyses through additional supplementary experiments.
6.1 Comparison on NLP Tasks
6.1.1 Datasets
As reported in the results above, we follow Yang et al. (2024b) and adopt five representative text classification datasets to evaluate merging methods under different protocols: AG News (4 classes) Zhang et al. (2015), Yelp (5 classes) yel (2014), Amazon (5 classes) McAuley and Leskovec (2013), Yahoo (10 classes) Zhang et al. (2015), and DBpedia (14 classes) Lehmann et al. (2015).
6.1.2 Implementation Details
For the backbone of models, $f_{b}$, we primarily use bert-base-uncased Devlin et al. (2018) as the base architecture, and the fine-tuned checkpoints are referenced from Yang et al. (2024b). Additionally, for a more comprehensive evaluation, we also adopt T5-base and T5-large Raffel et al. (2020) as $f_{b}$. For the scaling hyperparameters of the Task Arithmetic and Ties models, we follow the original paper by using a portion of labeled data for selection. These parameters are then fixed across all evaluation protocols to ensure a fair comparison. Our code is released in Github  [^1] https://github.com/fskong/FT-Classifier-for-Model-Merging.
6.1.3 Results Analysis
Each merging model is evaluated on all datasets using two protocols: Current Eval and FT-Classifier Eval, to calculate accuracy. To save space, we compute the average accuracy and present it using a line chart, with detailed results provided in the Appendix. To demonstrate that the FT-Classifier Eval maintains a certain level of robustness even with a small number of few-shot samples, we conduct experiments using five random seeds and report the mean and variance in the Figure.
The experimental results using BERT as $f_{b}$ have been reported in Figure 5 and Table 1. Unless otherwise specified, the experiments in this paper are based on BERT. And the results conducted on T5-base and T5-large are shown in Figure 6. From the Figure, we find that the main conclusions are essentially consistent with those on BERT, strengthening our arguments.
Refer to caption Refer to caption Refer to caption
(a) NLP tasks performance on T5-base.
Refer to caption Refer to caption Refer to caption
(b) NLP tasks performance on T5-large.
Figure 6: The performance of different evaluation protocols on NLP tasks under T5-base and T5-large. Detailed results are reported in Appendix A.3.
Refer to caption Refer to caption Refer to caption
(a) CV tasks performance on ViT-B-16.
Refer to caption Refer to caption Refer to caption
(b) CV tasks performance on ViT-B-32.
Figure 7: The performance of different evaluation protocols on CV tasks under ViT-B-16 and ViT-B-32. Detailed results are reported in Appendix A.4.
6.2 Comparison on CV Tasks
6.2.1 Datasets
For CV tasks, we also follow Yang et al. (2024b) and adopt eight representative image classification datasets: SUN397 (397 classes) Xiao et al. (2016), Cars (196 classes) Krause et al. (2013), RESISC45 (45 classes) Cheng et al. (2017), EuroSAT (10 classes) Helber et al. (2019), SVHN (10 classes) Netzer et al. (2011), GTSRB (43 classes) Stallkamp et al. (2011), MNIST (10 classes) Deng (2012), and DTD (47 classes) Cimpoi et al. (2013).
6.2.2 Implementation Details
For $f_{b}$, we primarily utilize ViT-B-16 and ViT-B-32 Radford et al. (2021) as the base architecture, and the fine-tuned checkpoints are referenced from Yang et al. (2024b). The selection for the merging scale is consistent with the description in the NLP tasks. It is worth noting that we found that, possibly due to the large number of classes in most CV datasets, using few-shot samples is insufficient to fine-tune the classifier effectively. Therefore, we increased the number of samples used for alignment learning. Specifically, we sampled subsets from the test set with sampling rates of {0.5%, 1%, 2.5%, 5%, 10%}.
6.2.3 Results Analysis
To verify that the misalignment issue also exists in CV tasks, we visualize the image embeddings using T-SNE. The results are shown in Appendix A.2. From the Appendix, we observe that the merging outputs also exhibit good clustering, indicating their high quality. Therefore, we believe that the misalignment problem also exists in CV tasks.
Further, to mitigate the misalignment issue in CV tasks, we applied the FT-Classifier protocol to re-evaluate different merging methods. The comparable results are shown in Figure 7. We find that with just over 1% of unlabeled data, the misalignment problem can be alleviated. This not only demonstrates the effectiveness of the FT-Classifier but also implicitly proves that misalignment is easy to capture.
6.3 Difference between $f_{m}$ and $f_{b}$
Previous analyses demonstrate that different $f_{m}$ achieves significant improvements when evaluated with a simple fine-tuning classifier. To further verify that the performance improvement stems from the inherent capability of the $f_{m}$ itself rather than additional information introduced by fine-tuning, we tested the results of the original pre-trained models $f_{b}$ under different evaluation protocols. The experimental results are presented in Table 3.
From the Table, we observe that regardless of the evaluation protocol, $f_{m}$ consistently outperforms $f_{b}$. This demonstrates that all $f_{m}$ models effectively incorporate knowledge from fine-tuned models. Meanwhile, our new evaluation protocols do not show significant improvements over the original protocol when applied to $f_{b}$, indicating that models with limited parameters cannot retain sufficient task-specific information. This proves that our evaluation protocols are essentially fair compared to the original ones and just activate the fine-tuned information that $f_{m}$ already contains.
Method $f_{b}$ WA TA Ties
Current Eval 17.5 55.8 64.3 65.1
FT-Classifier Eval 22.0 69.0 72.3 73.0
Table 3: Comparison of different $f_{m}$ and $f_{b}$ on different evaluation protocols.
7 Related Works
Model merging aims to combine multiple homologous models into a single model capable of performing better across multiple related tasks Wortsman et al. (2022); Jang et al. (2024); Daheim et al. (2023); Panigrahi et al. (2023); Marczak et al. (2024). Beyond basic Weight Averaging of multiple models, the primary merging methods can be categorized into two main types  Yadav et al. (2024).
Merging Weight Based Methods
Instead of assigning an averaging weight for each fine-tuned model, some methods utilize dataset statistics or gradients to adjust the merging weights. Fisher Merging Matena and Raffel (2022) computes per-weight importance for each model by the Fisher Information matrix, and reweights the weights with this importance factor during merging. RegMean Jin et al. (2022) conducts a local linear regression problem that minimizes prediction differences between the merged model and the individual models, and the closed-form solution is the merging weight. AdaMerging Yang et al. (2024c) autonomously learns task-wise or layer-wise merging coefficients without using the original training data. Specifically, it employs entropy minimization on unlabeled test samples as a surrogate objective to iteratively refine the merging coefficients.
Task Vector Based Methods
The task vector indicates the offset of a certain task based on the pre-trained model in the model parameter space. We introduce the following latest methods except for the representative referred to before. Motivated by model pruning that eliminating unnecessary parameters could also maintain performance, Dare Yu et al. (2024) reduces potential conflicts in the delta parameters by randomly dropping. To mitigate forgetting while training, MagMax Marczak et al. (2024) combines sequential fine-tuning with maximum-magnitude weight selection to enable effective knowledge integration across tasks. TALL-masks Wang et al. (2024) localizes task-specific information within the multi-task vector by deactivating irrelevant parts for each task using binary masks. Model Tailor Zhu et al. (2024) derives a sparse mask to identify the "model patch" using a fusion of salience and sensitivity analysis, then introduces a compensation mechanism to "decorate the patch", enhancing performance on both target and original tasks.
8 Conclusion
Model merging combines multiple fine-tuned models into a single one in the parameter space, enabling multi-task ability without full datasets or extra fine-tuning. We identify the misalignment issue between merging outputs and fine-tuned classifiers in classification tasks and show that addressing it improves classification performance. We also enhance evaluation protocols by fine-tuning an aligned classifier with few-shot examples, boosting evaluation quality with minimal time cost. Finally, we prove this misalignment can be converged as an orthogonal transformation, offering insights into model merging’s intrinsic properties.
Appendix A Appendix
A.1 Visualization of Sentence Embedding
T-SNE visualization of sentence embedding produced by different $f_{m}$ on Amazon, Yelp, and Yahoo are shown in Figure 8.
Refer to caption Refer to caption Refer to caption
(a) Weight Averaging (b) Task Arithmetic (c) Ties
Refer to caption Refer to caption Refer to caption
(d) Weight Averaging (e) Task Arithmetic (f) Ties
Refer to caption Refer to caption Refer to caption
(g) Weight Averaging (h) Task Arithmetic (i) Ties
Figure 8: T-SNE visualization of sentence embedding on Amazon, Yelp, and Yahoo.
A.2 Visualization of Image Embedding
The image embeddings produced by the $f_{m}$ model are shown in Figure 9 and 10. It is worth noting that, due to the large number of classes in the CV datasets, we did not display the T-SNE results of the fine-tuned model for better visualization. As shown in the Figure, we find that the embeddings produced by different $f_{m}$ outputs also exhibit high quality, which is consistent with the conclusions drawn from the NLP tasks.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 9: T-SNE visualization of image embedding for each merging method on CV tasks.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 10: T-SNE visualization of image embedding for each merging method on CV tasks.
A.3 Detailed Results on NLP Tasks
Detailed results on NLP tasks by T5-base and T5-large are presented in Table 4.
(a) T5-base
Method AG News Yelp Amazon Yahoo DBpedia Average
Fine-tuned 94.5 66.6 62.9 75.2 99.1 79.6
Weight Averaging 79.0 51.8 51.7 42.2 79.9 60.9
+FT-Classier 85.3 60.5 55.3 61.8 97.1 72.0
Task Arithmetic 80.6 59.2 55.2 51.5 79.5 65.2
+FT-Classier 88.2 64.0 58.2 67.6 98.4 75.3
Ties 80.0 59.2 55.4 50.7 79.3 64.9
+FT-Classier 88.3 63.8 58.1 68.0 98.4 75.4
(b) T5-large
Method AG News Yelp Amazon Yahoo DBpedia Average
Fine-tuned 94.9 68.6 64.9 77.1 99.1 80.9
Weight Averaging 79.2 55.5 44.3 49.0 94.5 64.5
+FT-Classier 88.0 62.6 59.7 67.9 98.2 75.3
Task Arithmetic 90.8 58.3 60.5 72.9 92.7 75.1
+FT-Classier 91.2 65.9 61.9 73.5 98.7 78.3
Ties 91.3 56.0 60.3 73.2 90.4 74.3
+FT-Classier 88.4 63.9 58.2 68.0 98.4 75.4
Table 4: Comparison on NLP tasks by T5. The results in this Table are based on a few-shot setting of $k=20$.
A.4 Detailed Results on CV Tasks
Detailed results on CV tasks by ViT-B-16 and ViT-B-32 are presented in Table 5.
(a) ViT-B-16
Method SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Average
Fine-tuned 81.8 86.8 96.9 99.7 97.8 99.1 99.7 82.0 92.9
Weight Averaging 67.7 70.0 75.3 79.5 74.9 60.1 94.4 43.8 70.7
+FT-Classier 69.7 72.7 83.8 89.7 73.7 84.3 95.3 51.5 77.6
Task Arithmetic 61.1 65.9 74.0 76.2 88.0 73.9 98.4 53.0 73.8
+FT-Classier 68.1 72.3 86.3 90.8 90.2 90.3 98.8 64.2 82.7
Ties 69.1 72.5 80.5 84.0 85.0 71.5 98.1 54.9 77.0
+FT-Classier 71.3 74.1 86.0 90.6 90.4 90.2 98.8 64.5 83.2
(b) ViT-B-32
Method SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Average
Fine-tuned 75.3 77.7 96.1 99.7 97.5 98.7 99.7 79.4 90.5
Weight Averaging 65.3 63.4 71.4 71.7 64.2 52.8 87.5 50.1 65.8
+FT-Classier 71.5 73.6 86.7 91.9 79.5 83.1 97.1 58.7 80.3
Task Arithmetic 55.2 54.9 66.7 78.9 80.2 69.7 97.3 50.4 69.1
+FT-Classier 62.5 61.5 80.4 89.9 86.4 89.9 98.4 61.3 78.8
Ties 65.0 64.4 74.8 77.4 81.2 69.3 96.5 54.5 72.9
+FT-Classier 73.1 77.0 89.1 93.9 88.0 89.7 98.7 68.1 84.7
Table 5: Comparison on CV tasks by ViT. The results in this Table are based on a sampling rate of 10%.