### v2

# Fine-tuning Aligned Classifiers for Merging Outputs: Towards a Superior Evaluation Protocol in Model Merging

Abstract
Model merging combines multiple fine-tuned models into a single one via parameter fusion, achieving improvements across many tasks. However, in the classification task, we find a misalignment issue between merging outputs and the fine-tuned classifier, which limits its effectiveness. In this paper, we first demonstrate the following observations: (1) Merging outputs exhibit the comparable cluster effect with fine-tuned outputs, and already contain necessary classification information; (2) The misalignment between merging outputs and the fine-tuned classifier can converge to an orthogonal transformation, and alleviating this misalignment can significantly enhance the performance of merging models. Based on these observations, we then propose a new protocol FT-Classifier, which fine-tunes an aligned classifier with few-shot unlabeled samples, enabling better evaluation of merging methods and improved classification performance.
Fine-tuning Aligned Classifiers for Merging Outputs:
Towards a Superior Evaluation Protocol in Model Merging
[…]
1 Introduction
Fine-tuning a pre-trained language model on downstream tasks to obtain a task-specific fine-tuned model is a common practice nowadays Devlin et al. (2018). Building on this training paradigm, real-world applications increasingly aim to leverage a single model that can be fine-tuned across multiple tasks, thus enabling it to address all tasks simultaneously Zhang and Yang (2021). This is known as multi-task learning (MTL), which generally aggregates all task data for joint training. However, MTL not only requires the full amount of task data, which raises data privacy concerns, but also involves a fully fine-tuning process, leading to high training costs and poor scalability Yang et al. (2024b). Therefore, researchers are actively exploring model merging as an alternative Izmailov et al. (2018); Jin et al. (2022); Ilharco et al. (2022); Yang et al. (2024c, a).
Model merging aims to directly combine multiple homologous fine-tuned task-specific models in the parameter space, without requiring the full dataset and the fine-tuning process, to obtain a merged model capable of solving multiple tasks simultaneously Yadav et al. (2023). Several studies have investigated model merging in areas such as continual learning Marczak et al. (2024), information retrieval Li et al. (2024), and classification Yu et al. (2024). Interestingly, we find that compared to other tasks, model merging brings fewer benefits in classification tasks.
We argue that this is due to the inevitable changes in the parameters of the merged model compared to the fine-tuned model. However, the current evaluation protocol (Current Eval) still utilizes the fine-tuned classifier for merged model assessment. This misalignment between merging outputs and the fine-tuned classifier leads to inaccurate evaluations of the merging method. In this work, we conduct the following analysis to validate our argument.
Merging outputs already exhibit a good clustering effect.
We visualize the merging outputs of three representative methods on text classification datasets using T-SNE Van der Maaten and Hinton (2008), as shown in Figure 1. From the Figure, we observe that although the performance of different merging methods varies across specific tasks, all merging outputs of representative methods exhibit comparable clustering performance to that of fine-tuned outputs. This motivates us to ponder that the merging outputs already include the necessary classification information, yet it is not accurately evaluated in the Current Eval.
Misalignment between merging outputs and fine-tuned classifier leads to the undervaluation of merging methods.
We bypass the fine-tuned classifier and instead use a few-shot set of labeled samples as class anchors. Employing a KNN-based method—commonly applied in sentence representation studies—we directly measure classification accuracy. The results of this KNN evaluation protocol (KNN Eval) are reported in Figure 2. Surprisingly, KNN Eval significantly outperforms the Current Eval on three representative methods, even with limited labeled data, demonstrating that Current Eval tends to underestimate merging methods. Since KNN Eval does not require a classifier at all, while Current Eval requires a fine-tuned classifier, this also proves that this underestimation may be caused by the misalignment between merging outputs and the fine-tuned classifier.
The misalignment can converge into an orthogonal transformation.
In conjunction with Figure 1, which illustrates a certain bias between merging outputs and fine-tuned outputs in the latent space, we hypothesize some simulation of transformation to align them. Specifically, we employ a series of transformations, varying the strength of the constraints, for empirical verification. The definition of different transformations and corresponding results are detailed in Table 1 and Figure 4, respectively. The results demonstrate that after transformation alignment, all merging methods exhibit significant improvement compared to the Current Eval, even approaching the upper bound of the fine-tuned model. Moreover, when these transformations converge to an orthogonal transformation, there is only a slight drop in classification performance. This indicates that through a rigid transformation such as rotation or reflection, the misalignment can be easily alleviated. As a result, the fine-tuned classifier may only need a simple alignment with the merging outputs to achieve better evaluation results.
Fine-tuning an aligned classifier can alleviate the misalignment and facilitate merging methods assessment.
Based on the above observations, we believe that Current Eval mainly assesses the compatibility between the merging outputs and the fine-tuned classifier rather than effectively evaluating the capability of the merging outputs themselves, which in turn limits the final classification performance. Therefore, we propose the FT-Classifier evaluation protocol (FT-Classifier Eval), which leverages few-shot unlabeled data to train an aligned classifier for merging outputs. This protocol does not introduce additional parameters and requires only minimal training steps, making it a convincing alternative to Current Eval. The results of the FT-Classifier Eval, as presented in Figure 5, demonstrate that such a lightweight and computationally feasible post-processing operation improves the effectiveness of merging methods assessment.
In short, we make the following contributions:
  • Through empirical analysis, we reveal that the misalignment issue between merging outputs and the fine-tuned classifier renders the Current Eval ineffective for accurately assessing merging methods.
  • We prove that this misalignment can converge to an orthogonal transformation, indicating that the intrinsic structure and geometric properties of the merging outputs remain consistent with those of the fine-tuned model.
  • We introduce a new FT-Classifier Eval that alleviates the misalignment issue using limited data, without adding parameters or changing the model structure. This enables a more accurate assessment of merging methods while significantly enhancing their performance.
2 Background
2.1 Definition of Model Merging
Let $f_{b}$ denote the pre-trained base model with parameter $\mathbf{\Theta}_{b}$. Given $T$ specific tasks with training data $X_{t}$ and corresponding label $Y_{t}$ for each data, one can obtain $T$ fine-tuned models. The fine-tuned model for task $t$ can be denoted as $f_{t}$ with parameter $\mathbf{\Theta}_{t}$. Without fully fine-tuning, model merging aims to merge these homologous models to generate a single $f_{m}$ with parameter $\mathbf{\Theta}_{m}$, which can simultaneously perform well across all $T$ tasks.
2.2 Representative Methods
Up to date, numerous model merging approaches have been proposed, which we will discuss in detail in Section 7. In this section, we primarily focus on introducing the three most representative methods.
Weight averaging Wortsman et al. (2022) simply averages $T$ fine-tuned models for merging, which is formulated as:
$\mathbf{\Theta}_{m}=\frac{1}{T}\sum_{t=1}^{T}\mathbf{\Theta}_{t}.$ (1)
Task Arithmetic Ilharco et al. (2022) defines the difference between the fine-tuned model and the pre-trained model in parameter space as a task vector, denoted by $\mathbf{\Delta}_{t}$. Experiments demonstrate that performing arithmetic operations on task vectors can effectively guide the merged model. The process is defined as:
$\mathbf{\Delta}_{t}$ $=\mathbf{\Theta}_{t}-\mathbf{\Theta}_{b},$ (2)
$\mathbf{\Theta}_{m}$ $=\mathbf{\Theta}_{b}+\lambda\sum_{t=1}^{T}\mathbf{\Delta}_{t},$
where $\lambda\in[0,1]$ is a scaling term determined by held-out validation sets.
Ties Yadav et al. (2023) posits that parameter conflicts are widespread across multiple task vectors. To address this, it preserves task parameters with large magnitudes and resolves sign conflicts using an election mechanism. We denote the task vector processed by Ties as $\mathbf{\Delta}_{t}^{\rm Ties}$, then the parameters merged by Ties are calculated as follows:
$\mathbf{\Theta}_{m}=\mathbf{\Theta}_{b}+\lambda\sum_{t=1}^{T}\mathbf{\Delta}_{t}^{\rm Ties}.$ (3)
2.3 Evaluation of Merging Models
Existing merging methods in classification tasks primarily focus on merging fine-tuned encoders, while not performing the merging operation on fine-tuned classifiers. That is, the merging encoders $f_{m}$ should be evaluated on the fine-tuned classier, which can be formulated as:
$\hat{Y}_{t}=f_{m}(X_{t};\mathbf{\Theta}_{m})^{\top}\mathbf{W}_{t},$ (4)
where $\mathbf{W}_{t}$ represents the weight of the fine-tuned classifier of task $t$, which maps the merging outputs to the predicted labels $\hat{Y}_{t}$. In the following descriptions, this currently adopted evaluation protocol is referred to as ‘Current Eval’.
3 Existence of the Misalignment
From Eq. 4, we can easily find the fact that due to the change in parameters, the merging outputs of $X_{t}$ generated by $f_{m}$ are inevitably shifted in the latent space compared to those generated by $f_{t}$. However, the evaluation of $f_{m}$ still relies on $\mathbf{W}_{t}$, which is tailored to $f_{t}$. Naturally, this raises a question: Does the outputs of $f_{m}$ align well with the fine-tuned classifier $\mathbf{W}_{t}$? To analyze this issue, this section delves into an in-depth analysis of the output of $f_{m}$.
3.1 Visualization of Merging outputs
Refer to caption Refer to caption
(a) Weight Averaging (79.2%) (b) Weight Averaging (55.1%)
Refer to caption Refer to caption
(c) Task Arithmetic (82.9%) (d) Task Arithmetic (81.5%)
Refer to caption Refer to caption
(e) Ties (83.3%) (f) Ties (83.4%)
Figure 1: T-SNE visualization of three representative merging methods on AG News and DBpedia. The blue points represent the sentence embeddings of the $f_{t}$ model, while the red points represent the results of the $f_{m}$ model. The color shades indicate different labels. The value indicates the classification accuracy of merging methods. For results on other datasets, please refer to Appendix A.1.
Motivated by Yang et al. (2024b), we used T-SNE to visualize the outputs of $f_{t}$ and three representative $f_{m}$ models. Furthermore, to observe the spatial variations of $f_{m}$’s output in more detail, we incorporated the corresponding label information into the visualization process, as shown in Figure 1. From this Figure, we observe two key phenomena:
(1) There is a representation bias between the merging outputs and the fine-tuned outputs. The overall red and blue clusters are non-overlapping, indicating that merging outputs and fine-tuned outputs exhibit a bias in the latent space. This phenomenon is consistent with the observations in Yang et al. (2024b).
(2) All merging outputs exhibit good clustering effects like fine-tuned outputs. Despite notable gaps in classification accuracy between different merging methods (e.g., nearly a $30\%$ accuracy difference between Weight Averaging and Ties on DBpedia), their merging outputs exhibit similar clustering performance. This suggests that all merging outputs retain necessary classification information, comparable to fine-tuned outputs.
Based on these two phenomena, we hypothesize that the misalignment between merging outputs and the fine-tuned classifier leads to an underestimation of merging methods in Current Eval.
3.2 KNN Evaluation of Merging outputs
To validate our hypothesis, we use a KNN-based method to directly evaluate the classification performance of different merging outputs. Specially, for the KNN evaluation (KNN Eval), we first randomly select few-shot labeled data as anchors for each class, then calculate the average Euclidean Distance between the test sample and these anchors for each class, and finally assign the class with the minimum distance as the prediction. The overall process is formulated as:
$\hat{y}_{t}=\arg\min_{y_{t}\in Y_{t}}\frac{1}{k}\sum_{i=1}^{k}d($ $f_{m}(X_{t};\mathbf{\Theta}_{m})),$ (5)
$f_{m}([A_{t}^{y_{t}}]^{i};\mathbf{\Theta}_{m}))).$
where $y_{t}$ is an element of $Y_{t}$, $k$ refers to the number of few-shot samples, $d$ represents the Euclidean Distance metric, $A_{t}^{y_{t}}$ represents the randomly selected anchors for label $y_{t}$, and []delimited-[][$\cdot$][ ⋅ ] denotes the indexing function. Illustration of Current Eval and KNN Eval are shown in Figure 3 (a) and (b).
Figure 2: The performance of KNN Eval compared with Current Eval under varying numbers of few-shot examples $k$. Fine-tuned represents the results obtained using $f_{t}$, which serves as the evaluation upper bound. The $x$-axis represents the $k$, and the $y$-axis represents the average Accuracy for all merging tasks. Corresponding detailed results while $k=5$ are shown in Table 2. Experimental details and explanations are provided in Section 6.
From Eq. 5, we observe that KNN Eval directly utilizes the merging outputs for classification with only a few labeled samples. We believe this evaluation protocol can effectively assess whether the necessary classification information is embedded in the merging outputs. Results of KNN Eval compared with Current Eval are shown in Figure 2.
From the Figure, we observe that KNN Eval surprisingly yields excellent classification results. Even with only 5 few-shot examples, the KNN Eval outperforms the Current Eval, which relies on a fine-tuned classifier. This demonstrates that the outputs of different merging methods contain sufficient information for classification. Therefore, the reason for the poor performance in the Current Eval is a misconception, more likely caused by the misalignment between the merging outputs and the fine-tuned classifier.
Figure 3: Illustrations of some variants referred to in this paper.
4 Analysis of the Misalignment
Through empirical analysis, we observe that there exists a misalignment issue between the merging outputs and the fine-tuned classifier, and this misalignment leads to the undervaluation of $f_{m}$’s capability in the Current Eval. Therefore, we hypothesize that: Alleviating the misalignment in latent space could lead to a better evaluation of $f_{m}$.
4.1 Different Transformation Hypothesis
To further explore this misalignment, we hypothesize a transformation between the merged outputs and the fine-tuned outputs, and we propose conjectures based on varying degrees of constraint on this transformation. Specifically, we align the merged outputs with the fine-tuned outputs by optimizing the following objective function:
$\arg\min_{G_{t}}||G_{t}(f_{m}(X_{t};\mathbf{\Theta}_{m}))-f_{t}(X_{t};\mathbf{\Theta}_{t})||_{1},$ (6)
where $G_{t}(\cdot)$ represents the transformation for task $t$, which can be instantiated using the transformations listed in Tabel 1. In particular, we employ unlabeled test data to train the related parameters. Noted, during the training process, only parameters in $G_{t}(\cdot)$ are updated, while all other parameters remain frozen. We refer to these merging outputs aligned method as ‘Aligned $f_{m}$’, illustrated in Figure 3 (c).
$G(\mathbf{x})=\mathbf{A}\mathbf{x}+\mathbf{b}$, where $\mathbf{A}\in\mathbb{R}^{n\times n},\mathbf{b}\in\mathbb{R}^{n}$
Transformation Constraints
Affine Transform /
Linear Transform $\mathbf{b}=0$
Invertible Linear Transform $\det(\mathbf{A})\neq 0,\mathbf{b}=0$
Euclidean Transform $\mathbf{A}^{\top}\mathbf{A}=\mathbf{I},\mathbf{b}\neq 0$
Orthogonal Transform $\mathbf{A}^{\top}\mathbf{A}=\mathbf{I},\mathbf{b}=0$
Scaling Transform $\mathbf{A}={\rm diag}(\mathbf{A}),\mathbf{b}=0$
Table 1: Definition of different transformations. When $\mathbf{x}$ is instantiated as the merging output, $G(\mathbf{x})$ represents the fine-tuned output. From Affine Transform to Scaling Transform, the constraints gradually become stronger.
Taking the orthogonal transformation as an example, we first randomly initialize a mapping matrix $\mathbf{A}_{t}$ and set $\mathbf{b}_{t}=0$ for each task $t$. Then, we perform optimization by imposing regularization constraints on the mapping matrix $\mathbf{A}_{t}$. The overall loss function is defined as follows:
$\arg\min_{\mathbf{A}_{t}}$ $||\mathbf{A}_{t}f_{m}(X_{t};\mathbf{\Theta}_{m})-f_{t}(X_{t};\mathbf{\Theta}_{t})||_{1}$ (7)
$+\alpha||\mathbf{A}_{t}^{\top}\mathbf{A}_{t}-\mathbf{I}||_{1},$
where $\alpha$ is the weight hyperparameter, and $\mathbf{I}$ is an identity matrix. By adjusting the value of $\alpha$, $\mathbf{A}_{t}$ can converge to an orthogonal transformation.
Results of all these Aligned $f_{m}$ are shown in Figure 4. From the Figure, we observe a very interesting phenomenon: When the transformation constraints are relaxed, the classification performance of all $f_{m}$ is barely affected and remains close to the upper bound (Fine-tuned results). For example, when the constraints are weaker than the orthogonal transformation, the performance is essentially unaffected, with at most nearly a 2% drop. This demonstrates that through an orthogonal transformation, the merging outputs can be aligned well with the fine-tuned outputs.
Figure 4: Comparison of different transformations. WA represents Weight Averaging, TA represents Task Arithmetic.
Method AG News Yelp Amazon Yahoo DBpedia Average
Fine-tuned model 91.6 60.4 57.9 72.7 98.8 76.3
Multi-task Learning 90.6 59.1 55.6 71.3 98.5 75.0
Weight Averaging 79.2 49.8 45.0 50.3 55.1 55.8
+ KNN Eval 81.9 $\pm$ 1.6 42.1 $\pm$ 4.3 41.2 $\pm$ 1.5 54.1 $\pm$ 3.5 91.5 $\pm$ 1.8 62.2 $\pm$ 0.6
+ FT-Classifier Eval 86.0 $\pm$ 1.0 51.5 $\pm$ 1.9 46.3 $\pm$ 0.6 66.2 $\pm$ 1.1 95.8 $\pm$ 0.4 69.2 $\pm$ 0.4
Task Arithmetic 82.9 55.8 48.4 53.1 81.5 64.3
+ KNN Eval 86.4 $\pm$ 1.3 54.2 $\pm$ 2.6 49.0 $\pm$ 1.6 61.6 $\pm$ 1.0 96.3 $\pm$ 0.2 69.5 $\pm$ 1.2
+ FT-Classifier Eval 87.5$\pm$ 0.7 55.6$\pm$ 0.5 51.5 $\pm$ 0.6 67.6 $\pm$ 0.7 96.6 $\pm$ 0.3 71.8 $\pm$ 0.2
Ties 83.3 55.7 47.0 55.9 83.4 65.1
+ KNN Eval 87.0 $\pm$ 1.6 54.6 $\pm$ 2.9 49.5 $\pm$ 1.5 63.4 $\pm$ 0.6 97.1 $\pm$ 0.2 70.3 $\pm$ 1.2
+ FT-Classifier Eval 88.1 $\pm$ 0.5 57.1 $\pm$ 0.9 52.6 $\pm$ 0.6 69.1 $\pm$ 0.6 97.7 $\pm$ 0.1 72.9 $\pm$ 0.1
Table 2: Comparison of FT-Classifier Eval and KNN Eval with the corresponding Current Eval (represented by merging method name). The results in this Table are based on a few-shot setting of $k=5$.
4.2 Further Discussion
Since orthogonal transformations in space are rigid, such as rotations or reflections, these transformations do not alter the intrinsic structure or geometric properties of the manifold. This demonstrates that the merging outputs largely retain the classification characteristics of the fine-tuned outputs. By further analyzing Figure 4 and considering the properties of orthogonal transformations, we can draw the following conclusions:
(1) With only a simple transformation with a limited number of parameters, the misalignment can be captured, achieving surprising classification performance that approaches the upper bound of the fine-tuned model. This further explains why merging outputs deliver strong performance even under direct classification (e.g., KNN Eval), as they preserve the essential information necessary for effective classification.
(2) The outputs of different $f_{m}$ contain comparable classification information, and Current Eval does not effectively assess $f_{m}$. Since the performance gap of different Aligned $f_{m}$ under the same transformation is much smaller than $f_{m}$. For example, the performance of Weight Averaging was initially around 10% lower than that of Task Arithmetic, but after alignment, the performance of Aligned Weight Averaging and Aligned Task Arithmetic becomes nearly identical.
These findings demonstrate that the misalignment issue hinders the Current Eval from effectively assessing the capacity of $f_{m}$. Clearly, alleviating the misalignment problem could potentially improve classification performance. Therefore, we hope to explore a new evaluation protocol that can effectively assess different merging methods by alleviating the misalignment.
5 FT-Classifier Evaluation Protocol
Figure 5: The performance of FT-Classifier Eval under varying numbers of few-shot examples $k$. Corresponding detailed results while $k=5$ are shown in Table 2.
Model merging combines the parameters of multiple fine-tuned models $f_{t}$ to produce a merged model $f_{m}$. Consequently, as discussed before, the output of $f_{m}$ inevitably shifts in the latent space relative to the original $f_{t}$, making this shift unavoidable. In this case, using $\mathbf{W}_{t}$, which is aligned well with $f_{t}$, to evaluate $f_{m}$ is inappropriate. The evaluation result reflects not the capability of $f_{m}$, but rather its compatibility with the fine-tuned classifier $\mathbf{W}_{t}$. Moreover, through experiments, we find that simply aligning the merging outputs and the fine-tuned outputs in the latent space can significantly improve the classification performance of $f_{m}$. Naturally, this motivates us to propose a new and reasonable evaluation protocol, which addresses the misalignment issue to more accurately assess $f_{m}$ and achieve improved performance.
Revisiting the objective of model merging, which aims to generate a new model $f_{m}$ using $f_{1},...,f_{T}$ that performs well across all $T$ tasks, we note that the upper bound of $f_{m}$’s performance on a specific task $t$ corresponds to the result from fine-tuned model $f_{t}$. Therefore, the objective of model merging is to minimize this performance discrepancy of $f_{m}$ and $f_{t}$ Ortiz-Jimenez et al. (2023). By optimizing this discrepancy, we could directly train an aligned classifier for $f_{m}$. That is, in the new evaluation protocol, we introduce an additional post-processing step while keeping the parameter scale and network structure unchanged, referred to as FT-Classifier Eval, which is defined as:
$\arg\min_{\mathbf{W}_{t}^{\prime}}\sum_{t=1}^{T}g(f_{m}(X_{t};\mathbf{\Theta}_{m})\!^{\top}\!\mathbf{W}_{t}^{\prime},f_{t}(X_{t};\mathbf{\Theta}_{t})\!^{\top}\!\mathbf{W}_{t})$ (8)
where $g$ is a measurement function. In this objective, we instantiate $g$ as KL divergence Kullback and Leibler (1951), and the new classifier $\mathbf{W}_{t}^{\prime}$ is initialized with the value of $\mathbf{W}_{t}$.
Refer to caption Refer to caption
(a) Performance on T5-base (b) Performance on T5-large
Figure 6: The performance of different evaluation protocols under T5-base and T5-large.
In model merging, obtaining few-shot data for each task is a relatively flexible requirement. Many studies use few-shot data to guide the merging methods Xiao et al. (2023), and some even rely on large-scale labeled or unlabeled data to determine merging weights Yang et al. (2024c); Jin et al. (2022) or the parameters to be pruned Ortiz-Jimenez et al. (2023); Zhu et al. (2024). Therefore, we also incorporate few-shot samples for aligned classifier training in FT-Classifier Eval.
Comparable results of Current Eval and FT-Classifier Eval are presented in Figure 5, and detailed results, including KNN Eval, utilizing 5-shot samples are shown Table 2. From the Figure, we observe that after simple fine-tuning even on 1-shot samples, FT-Classifier Eval performs significantly better than the Current Eval. This further demonstrates that the misalignment issue is highly manageable. Such a lightweight and computationally feasible post-processing operation facilitates the evaluation of $f_{m}$.
6 Experiments
6.1 Datasets
As reported in the results above, we follow Yang et al. (2024b) and adopt five representative text classification datasets to evaluate merging methods under different protocols: AG News (4 classes) Zhang et al. (2015), Yelp (5 classes) yel (2014), Amazon (5 classes) McAuley and Leskovec (2013), Yahoo (10 classes) Zhang et al. (2015), and DBpedia (14 classes) Lehmann et al. (2015).
6.2 Implementation Details
For the backbone of models, $f_{b}$, we primarily use bert-base-uncased Devlin et al. (2018) as the base architecture, and the fine-tuned checkpoints are referenced from Yang et al. (2024b). Additionally, for a more comprehensive evaluation, we also adopt T5-base and T5-large Raffel et al. (2020) as $f_{b}$. For the scaling hyperparameters of the Task Arithmetic and Ties models, we follow the original paper by using a portion of labeled data for selection. These parameters are then fixed across all evaluation protocols to ensure a fair comparison. Our code is released in Github  [^1] https://github.com/fskong/FT-Classifier-for-Model-Merging. For a more comprehensive analysis of the misalignment issue and the effectiveness of FT-Classifier Eval, we also introduce experiments on CV tasks, shown in Appendix A.2.
6.3 Results Analysis
Each merging model is evaluated on all datasets using three protocols: Current Eval, KNN Eval and FT-Classifier Eval. To save space, we compute the average accuracy and present it using a line chart. To demonstrate that the FT-Classifier Eval maintains a certain level of robustness even with a small number of few-shot samples, we use five random seeds and report the mean and variance.
The experimental results using BERT as $f_{b}$ have been reported in Figure 2, Figure 5 and Table 2. Unless otherwise specified, the experiments in this paper are based on BERT. All these results prove the existence of misalignment, and FT-Classifier Eval could alleviate this issue, achieving a better assessment of merging methods.
Results conducted on T5-base and T5-large are shown in Figure 6. From these results, we find that the main conclusions are essentially consistent with those on BERT, strengthening our arguments.
6.4 $f_{m}$ vs. $f_{b}$
Previous analyses demonstrate that different $f_{m}$ achieves significant improvements when evaluated with an aligned classifier. To further verify that the performance improvement stems from the inherent capability of the $f_{m}$ itself rather than additional information introduced by fine-tuning, we tested the results of the original pre-trained models $f_{b}$ under different evaluation protocols. The experimental results are presented in Table 3.
From the Table, we observe that regardless of the evaluation protocol, $f_{m}$ consistently outperforms $f_{b}$. This demonstrates that all $f_{m}$ models effectively incorporate knowledge from fine-tuned models. Meanwhile, FT-Classifer Eval does not show significant improvements over Current Eval when applied to $f_{b}$, indicating that models with limited parameters (only the classifier) cannot retain sufficient task-specific information. This proves that FT-Classifer Eval just activates the fine-tuned information that $f_{m}$ already contains.
Method $f_{b}$ WA TA Ties
Current Eval 17.5 55.8 64.3 65.1
KNN Eval 20.9 62.2 69.5 70.3
FT-Classifier Eval 22.0 69.2 71.8 72.9
Table 3: Comparison of different $f_{m}$ and $f_{b}$ on different evaluation protocols.
6.5 FT-Classifier vs. Surgery
The related work Surgery Yang et al. (2024b) characterizes the overall difference between the merging outputs and the fine-tuned outputs as “representation bias”, and reduces this bias by training additional parameters on a large amount of unlabeled data.
We argue that merging models naturally alters the output distribution (representation bias) but does not necessarily harm classification performance. The main cause of performance decline is the misalignment between the merging output and the fine-tuned classifier. We also show that this misalignment is simple (converging to an orthogonal transformation) and does not affect the intrinsic properties of merging outputs. Consequently, it can be mitigated by fine-tuning an aligned classifier for each task with very few samples. In addition, the FT-Classifier can also keep the parameter scale and network structure unchanged. In the context of model merging, we believe that such a lightweight and computationally efficient approach is more favorable. Results in Figure 7 demonstrate that the FT-Classifier outperforms Surgery when only a small number of samples are available.
7 Related Works
Model merging aims to combine multiple homologous models into a single model capable of performing better across multiple related tasks Wortsman et al. (2022); Jang et al. (2024); Daheim et al. (2023); Panigrahi et al. (2023); Marczak et al. (2024). Beyond basic Weight Averaging of multiple models, the primary merging methods can be categorized into two main types  Yadav et al. (2023).
Merging Weight Based Methods
Instead of assigning an averaging weight for each fine-tuned model, some methods utilize dataset statistics or gradients to adjust the merging weights. Fisher Merging Matena and Raffel (2022) computes per-weight importance for each model by the Fisher Information matrix, and reweights the weights with this importance factor during merging. RegMean Jin et al. (2022) conducts a local linear regression problem that minimizes prediction differences between the merged model and the individual models, and the closed-form solution is the merging weight. AdaMerging Yang et al. (2024c) autonomously learns task-wise or layer-wise merging coefficients without using the original training data. Specifically, it employs entropy minimization on unlabeled test samples as a surrogate objective to iteratively refine the merging coefficients.
Figure 7: Comparison of FT-Classifier and Surgery.
Task Vector Based Methods
The task vector indicates the offset of a certain task based on the pre-trained model in the model parameter space. We introduce the following latest methods except for the representative referred to before. Motivated by model pruning that eliminating unnecessary parameters could also maintain performance, Dare Yu et al. (2024) reduces potential conflicts in the delta parameters by randomly dropping. To mitigate forgetting while training, MagMax Marczak et al. (2024) combines sequential fine-tuning with maximum-magnitude weight selection to enable effective knowledge integration across tasks. TALL-masks Wang et al. (2024) localizes task-specific information within the multi-task vector by deactivating irrelevant parts for each task using binary masks. Model Tailor Zhu et al. (2024) derives a sparse mask to identify the “model patch” using a fusion of salience and sensitivity analysis, then introduces a compensation mechanism to “decorate the patch”, enhancing performance on both target and original tasks.
8 Conclusion
Model merging combines multiple fine-tuned models into a single one in the parameter space, enabling multi-task ability without full datasets or extra fine-tuning. In this paper, we first identify the misalignment issue between merging outputs and fine-tuned classifiers in classification tasks and show that merging outputs already contain necessary classification information. We then prove that this misalignment can be converged as an orthogonal transformation, offering insights into misalignment alleviation. We finally introduce an enhanced evaluation protocol, FT-Classifier Eval, by fine-tuning an aligned classifier with few-shot examples, boosting evaluation quality with minimal time cost.
Appendix A Appendix
A.1 Visualization of Sentence Embedding
T-SNE visualization of sentence embedding produced by different $f_{m}$ on Amazon, Yelp, and Yahoo are shown in Figure 8.
Refer to caption Refer to caption Refer to caption
(a) Weight Averaging (b) Task Arithmetic (c) Ties
Refer to caption Refer to caption Refer to caption
(d) Weight Averaging (e) Task Arithmetic (f) Ties
Refer to caption Refer to caption Refer to caption
(g) Weight Averaging (h) Task Arithmetic (i) Ties
Figure 8: T-SNE visualization of sentence embedding on Amazon, Yelp, and Yahoo.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 9: T-SNE visualization of image embedding for each merging method on CV tasks.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 10: T-SNE visualization of image embedding for each merging method on CV tasks.
Refer to caption Refer to caption
(a) CV tasks performance on ViT-B-16 (b) CV tasks performance on ViT-B-32
Figure 11: The performance of different evaluation protocols on CV tasks under ViT-B-16 and ViT-B-32.
A.2 Comparison on CV Tasks
A.2.1 Datasets
For CV tasks, we also follow Yang et al. (2024b) and adopt eight representative image classification datasets: SUN397 (397 classes) Xiao et al. (2016), Cars (196 classes) Krause et al. (2013), RESISC45 (45 classes) Cheng et al. (2017), EuroSAT (10 classes) Helber et al. (2019), SVHN (10 classes) Netzer et al. (2011), GTSRB (43 classes) Stallkamp et al. (2011), MNIST (10 classes) Deng (2012), and DTD (47 classes) Cimpoi et al. (2014).
A.2.2 Implementation Details
For $f_{b}$, we primarily utilize ViT-B-16 and ViT-B-32 Radford et al. (2021) as the base architecture, and the fine-tuned checkpoints are referenced from Yang et al. (2024b). The selection for the merging scale is consistent with the description in the NLP tasks. It is worth noting that we found that, possibly due to the large number of classes in most CV datasets, using few-shot samples is insufficient to fine-tune the classifier effectively. Therefore, we increased the number of samples used for misalignment alleviation. Specifically, we sampled subsets from the test set with sampling rates of {0.5%, 1%, 2.5%, 5%, 10%}.
In the ViT series of experiments, the image input is processed through the image encoder, and text inputs (label names) are processed through the text encoder, obtaining the image embedding $\mathbf{E}_{image}$ and text embeddings $\mathbf{E}_{text}$ respectively. Classification is performed by matching $\mathbf{E}_{image}$ with $\mathbf{E}_{text}$. In the settings of most model merging papers, the text encoder is frozen, meaning that $\mathbf{E}_{text}$ remains unchanged. In this case, we consider $\mathbf{E}_{text}$ can be understood as the classifier parameter $\mathbf{W}_{t}$, that is, initializing the classifier with label name embeddings. This approach is consistent with the way many text classification methods initialize label embeddings.
A.2.3 Visualization of Image Embedding
The image embeddings produced by the $f_{m}$ model are shown in Figure 9 and 10. It is worth noting that, due to the large number of classes in the CV datasets, we did not display the T-SNE results of the fine-tuned model for better visualization. As shown in these Figures, we find that the embeddings produced by different $f_{m}$ also exhibit good clustering performance, which is consistent with the conclusions drawn from the NLP tasks. Therefore, we believe that the misalignment problem also exists in CV tasks.
A.2.4 Results Analysis
Further, to mitigate the misalignment issue in CV tasks, we applied the KNN Eval and FT-Classifier Eval to re-evaluate different merging methods. The comparable results are shown in Figure 11. We find that with just over 1% of unlabeled data, the misalignment problem can be alleviated. This not only demonstrates the effectiveness of the FT-Classifier Eval but also implicitly proves that misalignment is easy to capture.